### v1

Abstract
GPU-enhanced architectures are now dominant in HPC systems, but message-passing communication involving GPUs with MPI has proven to be both complex and expensive, motivating new approaches that lower such costs. We compare and contrast stream/graph- and kernel-triggered MPI communication abstractions, whose principal purpose is to enhance the performance of communication when GPU kernels create or consume data for transfer through MPI operations. Researchers and practitioners have proposed multiple potential APIs for stream and/or kernel triggering that span various GPU architectures and approaches, including MPI-4 partitioned point-to-point communication, stream communicators, and explicit MPI stream/queue objects. Designs breaking backward compatibility with MPI are duly noted. Some of these strengthen or weaken the semantics of MPI operations. A key contribution of this paper is to promote community convergence toward a stream- and/or kernel-triggering abstraction by highlighting the common and differing goals and contributions of existing abstractions. We describe the design space in which these abstractions reside, their implicit or explicit use of stream and other non-MPI abstractions, their relationship to partitioned and persistent operations, and discuss their potential for added performance, how usable these abstractions are, and where functional and/or semantic gaps exist. Finally, we provide a taxonomy for stream- and kernel-triggered abstractions, including disambiguation of similar semantic terms, and consider directions for future standardization in MPI-5.
1 Introduction
Most MPI implementations now include support for GPU data paths, enabling standard MPI host-based APIs to move data to and from GPU memory. This feature is generally termed GPU-aware MPI and significant progress has been made toward standardizing this implementation feature through work on MPI memory types [11]. Importantly, the MPI operation stages of initiation, starting, completion, and freeing [10] are all conducted by code running on the host CPU in this approach; GPU code that needs to cause or coordinate MPI operation state changes must do so indirectly by synchronizing with CPU code that makes those updates on behalf of the GPU code.
Integrating MPI into GPU control paths is also important because it could eliminate the GPU/host communication latencies in current MPI implementations. This is a more challenging problem since it requires addressing multiple complex issues, including:
  • Which GPU control paths (stream execution and kernel execution) should support interacting with MPI operations?
  • To what degree can current MPI abstractions be reused on GPU control paths?
  • What are the concurrency semantics of MPI operations on host and GPU control paths?
The myriad proposed approaches for addressing these issues [1, 3, 5, 9, 10, 12, 13, 15, 16] have focused on GPU-triggered communication APIs in which host calls initialize and free MPI communication requests and the GPU control path initiates and in some cases completes MPI communication requests. When the GPU control path used for initiation and completion is the GPU stream scheduler, here and elsewhere this is termed stream-triggered communication and generally does not involve GPU-callable routines; instead the host API interacts with the GPU stream control program to schedule implicit GPU initiation and completion of communication operations. In contrast, using GPU kernel thread execution for initiation and completion is generally termed kernel-triggered communication and requires the addition of explicit GPU-callable APIs for initiating and completing communication.
Despite the many proposals, there is neither a consensus nor standard, either formal or informal, on an appropriate MPI API for GPU-triggered communication. To provide information and guidance for the MPI community to systematically address this issue (such as in MPI-5, or ideally sooner), this paper collects, organizes, and summarizes information about the various proposed MPI GPU-triggered communication approaches and highlights their similarities and differences. After providing background on representative GPU communication triggering proposals (Section 2), this paper makes the following contributions:
  • Presents a taxonomy for approaches to GPU-triggered communication that highlights the major design decisions that an MPI API needs to address (Section 3);
  • Summarizes the key features of and classifies eight different MPI stream- and kernel-triggering API proposals using this taxonomy and compares and contrasts key requirements and API design elements of these approaches (Section 4);
  • Highlights key gaps in the existing GPU triggering MPI APIs and in the MPI standard that will need to be addressed by future proposals for a robust MPI GPU triggering API (Section 5).
Finally, the paper offers conclusions and recommends a path forward in Section 6.
2 Representative GPU-Triggered Communication Abstractions
To illustrate the wide range of design choices and challenges in this area, we begin by providing information on three proposals for GPU-triggered MPI communication. Specifically, we discuss the stream triggered, two-sided communication support in recent MPICH implementations [16], the stream-triggered MPI one-sided communication support recently proposed by researchers at Hewlett Packard Enterprise (HPE) [13], and extensions to the MPI 4.0 Partitioned Communication interface [10] to support GPU kernel triggering of MPI communication [8]. We also present simple ping-pong communication for each proposal to further elucidate their proposed approaches. Together, this section highlights and provides essential background on the diverse design choices we explored in designing the GPU triggered communication taxonomy described below in Section 3. Note that while the examples in this section utilize CUDA APIs, appropriate ROCM APIs could be utilized instead.
2.1 MPICH Triggering
The GPU triggering support in recent MPICH implementations [16] focuses onstream triggering of MPI communication using two abstractions:
MPIX_Stream
–an MPI object to sequence MPI operations and to which external execution contexts, for example GPU streams, can be bound. This provides clearly defined concurrency semantics for MPI operations in multiple execution contexts including multi-threaded host execution contexts and accelerator-based MPI+X environments; and,
GPU Stream communicators
–MPI communicators to which an MPIX_-
Stream
object is attached to a GPU stream execution context, and on which MPI operations are enqueued to the attached GPU stream and return immediately in the host execution context.
MPICH provides explicit enqueue variants of the supported GPU-triggered operations that return immediately to the host; GPU synchronization routines must be used to await for completion of these operations on the GPU. MPICH implementers considered making MPI_Send and MPI_Recv on communicators with GPU streams attached have enqueue semantics (similar to the previously-proposed MPI-GDS API [15], but in the end rejected this because it would change the semantics of well-known MPI APIs. Finally, MPICH stream communicators also include additional features to support APIs similar to the previous MPI Endpoints proposal [2].
1 MPIX_Stream mpistream;
2 MPI_Comm stream_comm;
3 MPI_Info info;
4
5 MPI_Info_create(&info);
6 MPI_Info_set(info, "type", "cudaStream_t");
7 MPIX_Info_set_hex(info, "value", &cudastream,
8 sizeof(cudastream));
9 MPIX_Stream_create(info, &mpistream);
10 MPI_Info_free(&info);
11 MPIX_Stream_comm_create(MPI_COMM_WORLD, mpistream,
12 &stream_comm);
13
14 for (int i = 0; i < niters; i++) {
15 if (my_rank == 0) {
16 MPIX_Send_enqueue(src_buf, 1, MPI_INT, 1, 123,
17 stream_comm);
18 MPIX_Recv_enqueue(src_buf, 1, MPI_INT, 1, 123,
19 stream_comm, MPI_STATUS_IGNORE);
20 } else if (my_rank == 1) {
21 MPIX_Recv_enqueue(dst_buf, 1, MPI_INT, 0, 123,
22 stream_comm, MPI_STATUS_IGNORE);
23 MPIX_Send_enqueue(dst_buf, 1, MPI_INT, 0, 123,
24 stream_comm);
25 }
26 }
27 cudaStreamSynchronize(cudastream);
Figure 1: Stream-triggered MPI Ping Pong using MPICH Stream Triggering [16]
Figure 1 shows an example ping-pong communication using this interface. We highlight the following key features of this example:
  • To use GPU stream-triggered communications, the programmer must create a new communicator;
  • The stream communicator concept, similar to the one previously proposed for MVAPICH [15], can be naturally extended to include additional operations beyond the point-to-point operations supported in the current implementation, including collective communications; and,
  • MPICH’s stream communicators enqueued operations require the calling process must explicitly synchronize with the underlying GPU stream using GPU-specific operations to ensure completion of the corresponding MPI operation in the GPU execution context.
2.2 HPE One-Sided
The HPE one-sided GPU triggering API [13] supports stream-triggered, one-sided communication [^1] HPE also provides stream-triggered, two-sided communication, not considered here.. The stream abstractions included in HPE’s approach rely on GPU-NIC asynchronous stream triggering that integrates the NIC command queue and the GPU stream. In this approach, the CPU puts commands in the Cray Cassini NIC’s command queue and then GPU stream operations triggers these commands. This removes unnecessary synchronization on the CPU. HPE’s one-sided API is designed to allow for GPU-NIC progress where the host CPU is fully bypassed in progressing communication. HPE’s previous two-sided GPU triggering APIs [12] required host progress to receive messages due to the need for tag magtching and unexpected message handling.
1 /* Normal MPI communicator, window, group, and */
2 /* buffers assumed to exist */
3 for(int i = 0; i < niters; i++) {
4 if(rank == 0){
5 /* Send ping */
6 MPI_Win_start(group, MPI_MODE_STREAM, win);
7 MPI_Put(src,n,MPI_INT,1, disp, n, MPI_INT,win);
8 /* Puts triggered here */
9 MPIX_Win_complete_stream(win,stream);
10 /* Receive pong */
11 MPIX_Win_post_stream(group, win, stream);
12 MPIX_Win_wait_stream(win,stream);
13 } else { /* Receive ping */
14 MPIX_Win_post_stream(group, win, stream);
15 MPIX_Win_wait_stream(win, stream);
16 /* Send pong */
17 MPI_Win_start(group, MPI_MODE_STREAM, win);
18 MPI_Put(src, n, MPI_INT, 0, disp, n, MPI_INT,win);
19 /* Puts triggered here */
20 MPIX_Win_complete_stream(win, stream);
21 }
22 }
23 cudaStreamSynchronize(stream);
Figure 2: Stream-triggered MPI Ping Pong using HPE One-Sided Stream Triggering
Figure 2, shows an example of a one-sided ping-pong test using the HPE Cray MPI interface where stream-triggered variants of window post, wait, and complete operations are combined with window start and put operations to achieve fully offloaded GPU communication. We highlight the following key features of this example:
  • One-sided communication mixes new MPIX_* functions and standard MPI functions, because the Window is associated with the stream.
  • MPI_Put and MPI_Win_start do not have stream-triggered MPIX_ alternatives and simply initialize future communications operations to be triggered by the stream-triggered complete operation.
  • The MPI_Win_complete_enqueue initiates operations initialized by the host in the command queue. This utilizes the GPU and NIC without weakening MPI RMA synchronization semantics, and in reality strengthens MPI RMA semantics by always deferring one-sided operations until MPI_Win_complete_ enqueue is executed by the GPU stream control program.
2.3 Kernel-Triggered Partitioned Communication
1__global__ void ready(MPIX_Prequest preq) {
2 MPIX_Pready(threadIdx.x, preq);
3}
4
5__global__ void arrived(MPIX_Prequest preq) {
6 int flag = -1;
7 for(;;){
8 MPIX_Parrived(preq, threadIdx.x, &flag);
9 if(flag) break;
10 }
11}
12
13void pingpong(...) {
14 int sendbuf[1], recvbuf[1];
15 MPIX_Request srequest, rrequest;
16 MPIX_Prequest psrequest, prrequest;
17 int otherrank = (rank == 0) ? 1 : 0;
18 MPIX_Psend_init(sendbuf, 1, 1, MPI_INT, otherrank, 0,
19 MPI_COMM_WORLD, MPI_INFO_NULL,
20 &srequest);
21 MPIX_Precv_init(recvbuf, 1, 1, MPI_INT, otherrank, 0,
22 MPI_COMM_WORLD, MPI_INFO_NULL,
23 &rrequest);
24 MPIX_Prequest_create(srequest, &psrequest);
25 MPIX_Prequest_create(rrequest, &prrequest);
26 for(int i = 0; i < maxiters; i++) {
27 MPIX_Start(&rrequest); // ping if rank==1, else pong
28 MPIX_Start(&srequest); // pong if rank==1, else ping
29 if(rank == 0) {
30 ready<<<1, 1, 0, 0>>>(psrequest);
31 MPIX_Wait(&srequest, MPI_STATUS_IGNORE);
32 arrived<<<1, 1, 0, 0>>>(prrequest);
33 MPIX_Wait(&rrequest, MPI_STATUS_IGNORE);
34 } else {
35 arrived<<<1, 1, 0, 0>>>(prrequest);
36 MPIX_Wait(&rrequest, MPI_STATUS_IGNORE);
37 ready<<<1, 1, 0, 0>>>(psrequest);
38 MPIX_Wait(&srequest, MPI_STATUS_IGNORE);
39 }
40 }
41}
Figure 3: MPI Ping Pong using MPI-ACX Partitioned Stream Triggering
The MPI-ACX implementation [1] includes a prototype of the partitioned communication detailed in MPI-4. Partitioned point-to-point communication was specifically added to MPI-4 for a variety of reasons, including providing matching semantics (matching at initialization) and lightweight interfaces for data transfer initiation  [^2] It also removes the restrictions of persistent send and receive operations—MPI_Send_init and MPI_Recv_init—from MPI-1, which are forbidden to communicate and do not match only once. These fail to offer a channel semantic or obviate receive-queue matching. By contrast, single partitioned mode provides a uni-directional channel concept.. A specific nuance in this approach is that both MPIX_ Parrived and MPIX_Pready can be called from the host and the CUDA device. There are device-variants of the host requests to support the device-invoked operations.
Notably, the absence of a strong progress guarantee in MPI limits the guarantee that the partitioned communication initialization phase (i.e., MPI_Psend_init has already been matched proactively with MPI_Precv_init) when MPI_Start is called; it can instead occur at the first MPI_Wait for both sides of the partitioned channel. As such, proposals were offered for MPI-4.1 denoted MPI_Pbuf_prepare to guarantee initialization completion prior to first use and ensure readiness of buffers upon reuse [6]. This has not yet been adopted for the standard, but appears essential to avoid unexpected messages that would introduce complex logic that is considered untenable in GPU kernels.
Figure 3 details a ping-pong test based on the ring program supplied with MPI-ACX; and the following are some key observations:
  • In MPI-ACX, the programmer uses MPIX_Request to create a MPIX_Prequest for the device to use in operations. These objects are tied to each other with MPIX_Prequest_create.
  • The host-based MPIX_Request object is used for regular MPI functions such as MPIX_Start and MPIX_Wait.
  • Different request types and funcation signatures are needed for GPU and CPU interfaces due to differences between the state needed in each context.
3 A Taxonomy of MPI GPU-Triggered Communication
Based on a careful examination of both the MPI standard and prior proposals for GPU-triggered communication, we created the following categories for classifying key characteristics of MPI GPU-triggered communication proposals in four different overarching areas:
Area 1: GPU control path used:
Does the API support the GPU stream control path or kernel control path for initiating and completing MPI communication operations? In the case of APIs where both stream and kernel triggering are supported, we split the API into separate stream- and kernel-triggered sub-APIs and classify them separately because of the frequent semantic and API differences between the such approaches in other categories.
Area 2: API Design Considerations
Reuses Existing MPI APIs or abstractions:
Does the GPU triggering API reuse existing MPI communication operations or abstractions to enable GPU triggering of communication operation initiation or completion, or does it instead create new MPI operations or abstractions?
Changes Existing MPI API Semantics:
If the API reuses existing APIs or abstractions, does it change the semantics of these API elements to behave differently than they currently do? For proposals that modify the API, we classify whether the API strengthens MPI API semantics in a way that preserves backward compatibility or weakens MPI semantics in a way that may compromise backward compatibility.
Separate MPI Operation Initialization and Initiation:
Does the API use a single call to execute the starting stages of MPI communication operations or does it separate this into explicit API calls to initialize and initiate MPI communication operations?
GPU MPI Operation Completion Support:
Does the API support GPU detection of the completion of GPU-initiated MPI operations? We identify the classifications of all, some, and none for this category, corresponding to GPU control paths supporting waiting or polling for the completion of (1) all GPU-initiated MPI operation, (2) some GPU-initiated MPI operations, or (3) requiring host code to wait or poll for the completion of MPI operations.
Collective communication support:
Does the API specify operations enabling a single GPU communication operation to communicate with multiple processes? We classify collective support as either full, group, or none, cooresponding to (1) an API that supports the full semantics of MPI two-sided collecgtives, (2) an API that supports collective pairwise exchanges between subgroups of a communicator such as one-sided group or two-sided neighbor collectives, or (3) no specified support for collective communication.
Area 3: Ordering and Concurrency Considerations
MPI operation sequencing abstraction:
Does the API include an explicit MPI abstraction or object for ordering GPU communication operations separate from GPU-provided abstractions?
Sequencing Abstraction semantics:
If the API includes an explicit sequencing abstraction, are communication operations sequenced using that abstraction fully ordered or partially ordered (e.g., using a fencing or graph abstraction)?
Specification of MPI Concurrency Standard Integration:
Does the API address concurrency-related issues arising from GPU triggering of MPI communication through explicit specification or implicit specification via the existing MPI thread multiple mode, or are concurrency semantics unspecified?
Area 4: Implementation Considerations
GPU/NIC Progress:
Does the API enable message initiation and completion by the GPU and network device without the involvement of the host CPUS? This is generally accomplished by supporting one-sided data movement that avoids the control flow complexity of matching in the receive path.
Available Implementation or Evaluation:
Is an implementation or evaluatio of this API available either for download and installation, from a vendor, or in a peer-reviewed publication so that it can potentially be tested or compared with other APIs?
Multi-architecture Support:
Does the implementation, if available, target a single vendor architecture or has it been demonstrated for multiple GPU and/or NIC accelerator architectures? We categorize an implementation as Yes when it supports multiple GPUs and network architectures, as GPU or NIC when it has been demonstrated on multiple GPU or NIC architectures, respectively, and as No when it has been demonstrated on neither multiple NIC nor multiple GPU architectures.
4 Classifying Existing MPI GPU Triggering Proposals
In this section, we classify these proposals and their availablity or described implementations according to the taxonomy described in Section 3, as shown in Table 1. We then use this taxonomical classification to compare and contrast key features of these proposals and their corresponding implementations.
4.1 API Classification
We focus on the eight proposed GPU-triggered MPI programming interface extensions listed below:
MPI-GDS [15]
– a GPU-triggering interface extension for the MVAPICH MPI implementation [14].
MCI-ACX Enqueued [1]
– stream-triggered two-sided communication operations as implemented in the MPI-ACX GPU-triggering interface extension for NVIDIA network devices that support MLX4 device functionality.
MPICH Triggering [16]
– a GPU triggering interface extension for the MPICH MPI implementation [7].
HPE Send/Recv [12]
– a stream-triggered MPI interface extension for HPE systems with the Slingshot network interface [4].
Project Delorean [9]
– a proposal for graph-based sequencing of MPI operations that including GPU triggering.
HPE One-sided [13]
– a GPU-triggered Post/Start/Complete/Wait interface extension for HPE systems with the Slingshot network interface.
Partitioned Communication
– the MPI partitioned communication API [10] supplemented with the Pbuf_prepare and partitioned collective operations recently proposed [8], and partially implemented by MPI-ACX [1].
HPE Persistent [5]
– a kernel-triggered communication MPI interface extension for HPE systems with the Slingshot network interface [4].
Note that multiple variants of some of these designs have been proposed in the MPI Forum Working groups, BOFs, and other venues; we focus on the above listed API descriptions and implementations as representative of the main existing approaches at the time of this publication.
Table 1: Taxonomy of MPI GPU-Triggered Communication Proposals
Area 1 Area 2: API
Features
Proposal Control Reuses Changes Separate GPU Collective
Path Existing API Initialize Completion Support
Used APIs Semantics and Initiate Support
MPI-GDS Stream Yes Weaker No Full Full1
MPI-ACX Stream Yes1 No Yes1 Full No
Enqueued
MPICH Stream Yes No2 No Full No
Triggering
HPE Stream No No Yes Full No
Send-Recv
Delorean Stream No3 No Yes Full Full
HPE Stream Yes Stronger4 Yes Full Group
One-sided
Partitioned Kernel Yes No Yes Partial5 Full1
Comm.
HPE Kernel Yes No Yes Full No
Persistent
Area 3: Ordering and Area 4: Implementation
Concurrency Considerations
Proposal MPI Sequencing Concurrency GPU/NIC Impl. Multiple
Sequencing Model Model Progress Available Arch.
Object Semantics Specified Support
MPI-GDS No N/A Implicit No Yes No
MPI-ACX No N/A No No Yes No
Enqueued
MPICH Yes Full Explicit No Yes Yes
Triggering
HPE Yes Partial No No Yes GPU
Send/Recv
Delorean Yes Partial Implicit No No No
HPE No N/A Implicit Yes Yes GPU
One-sided
Partitioned No N/A Explicit Yes Yes No
Comm.
HPE No N/A No No Yes GPU
Persistent
  • 1
    Proposed in publication but not in the described or available implementations.
  • 2
    The initial proposal considered changing the semantics of MPI_Send and MPI_Recv on GPU stream communicators to be identical to their _enqueue variants and return to the caller prior to local completion of the operation; MPICH developers have since abandoned this approach.
  • 3
    Proposes “deferred” operations as a generalization of MPI persistent operations.
  • 4
    Explicitly defers MPI_Put/MPI_Get execution to when MPI_Win_complete_stream is called.
  • 5
    There is no GPU call to determine if a partition that has been marked ready to send has completed locally.
4.2 API Feature Comparisons
Based on the comparison shown in Table 1, there is no clear agreement on whether stream or kernel triggering is preferable, either from an API, programmability/usability, implementation, or performance perspective. Most implementations seek to reuse portions of the API and maintain or strengthen the semantics of the API operations they use, and also separate initialization and initiation of API operations.
Almost every API allows the GPU to check for or wait for the completion of every MPI operation that they initiate, allowing the GPU to avoid returning to the host execution context to ensure communication completion. The main exception is partitioned communication, where receives completion can be tested on the GPU using Parrived, but there is no equivalent call for testing for the completion of sends initiated from the kernel using Pready (the only available option is to check if all send partitions have completed using a single call to Wait or Test or one of their variants).
While several APIs propose collective support, only limited support for any GPU-triggered collective communication is available in the form of point-to-point group communication in the HPE one-sided triggering API. This is likely due to the complexity of combining complex collective communication with the constrained GPU and NIC control paths.
There is little agreement on whether to introduce a new MPI-level primitive to represent GPU execution ordering, with some APIs introducing new primitives and others reusing opaque GPU streams. Note that this is the case even from the same vendor, where the HPE two-sided API includes an MPI_Queue object that is not used by the HPE one-sided API.
Implementation availability also varies significantly. For instance, Delorean is just a specification with no known implementation; MPI-ACX is fully available and we were able to run the ping-ping example on representative hardware. Few implementations support full GPU/NIC offload of communication, with most APIs designed with the assumption that the host will be involved in data movement.
5 Key Gaps
Based on the classifications described in the previous section, this section highlights covers gaps and perceived limitations observed in prior work and prototypes as well as the MPI standard.
5.1 Limitations in the MPI Standard
In the MPI Standard, we observe that there are two main areas that needs to be addressed.
  1. 1.
    Problems surrounding persistence:
    The separation of initialization and initiation in GPU triggering APIs make persistent operations attractive for this purpose, as demonstrated by partitioned communication. Unfortunately, MPI-1 specified persistent send and receive operations cannot communicate and their initialization does not result in matching; indeed, persistent and non-persistent operations can match each time they are initiated.
  2. 2.
    Lack of MPI_Pbuf_prepare and similar operations for two-sided communication:
    There are two separate issues that have been conflated in the prior discussions of MPI_Pbuf_prepare and MPI_Pbuf_prepareall, yet to be standardized [6]:
    1. (a)
      Ensuring that MPI two-sided requests are matched (once and for all) so that there is agreement between the sender and receiver side about the exposed buffer on which to make operations (such as with MPI_Puts for early-bird communication as MPI_Pready’s are issued on the send side).
    2. (b)
      Ensuring or asserting that the remote buffer is ready for one-sided data movement; that is, there are no unexpected transfers.
    The former evidently subsumes the latter; otherwise, matching/synchronization would have to be done prior to every call to MPI_Start, which is likely to be unacceptably high.
    Further, because of the allowance for weak progress in MPI implementations, there is no guarantee even in two sided operations that match at initialization (e.g. partitioned communication) that the initial match has occured between before the first complete communication. This means that initial communication can happens before any agreement on the readiness of the receive buffer; for instance, agreement and transmission could first happen during the first wait calls, or the initial transfers could be done differently than subsequent ones.
    In addition, the desire for these conditions is not unique to partitioned communication — the various stream-triggered communication proposals (MPICH, HPE two-sided, MPI-ACX enqueuing, etc.) would want this functionality to push their data movement through the one-sided path as well.
5.2 Limitations across Proposed APIs
In addition to gaps in the MPI standard that increase the difficulty of supporting GPU triggering, there are also recurring limitations across the various APIs that have been proposed. We highlight these gaps below to encourage efforts to address them consistently in new proposals:
  • Clear concurrency model—If the host execution context and accelerator context both initiate operations, the ordering and concurrency is ill-defined in many APIS; any (left to the vagueness of MPI), or are silent. MPICH stream triggering is the only one that provides a clear, specific concurrent model for both host- and accelerator-initiated communication.
  • Matching—There is no explicit support for matching of send and receive buffers so that this occurs only once, and receive queue matching overheads are eliminated.
  • GPU completion support with partitioning—If something is initiated on a GPU, there is no way to know if something is completed on the send side. This makes it difficult or impossible to write programs that do not need to synchronizae with the CPU to interleave communication and computation.
  • Collective communication support—This is broadly absent and needed to support real applications that avoid unnecessary synchronization with the CPU. In particular, no current stream triggering approach supports full stream-triggered collective communication.
  • GPU/NIC Progress—Lack of this feature is a gap for any performant implementation. Out of the eight APIs considered only two (partitioned communication augmented with Pbuf_prepare and HPE one-sided communication) support GPU/NIC progress.
6 Conclusion
This is the first comprehensive survey of all eight proposals known to the authors that address stream/graph- and kernel-triggered communication abstractions for MPI, whose principal purpose is to improve the performance of communication when GPU kernels create or consume data for transfer through MPI operations. As stated at the outset, GPU-enhanced architectures are now dominant in HPC systems, but message-passing communication involving GPUs with MPI has proven to be both complex and expensive, motivating new approaches that lower such costs.
Here, we compared and contrasted stream/graph- and kernel-triggered MPI communication abstractions. Various proposals with distinct potential APIs for stream- and/or kernel-triggering have been proffered as designs and/or prototypes, which span various GPU architectures and approaches, including MPI-4 partitioned point-to-point communication, stream communicators, and explicit MPI stream/queue objects. Notably, some of these proposals either strengthen or weaken the semantics of MPI operations. We enumerated eight different proposals and provided a taxonomy of these proposals and their operations against twelve categories of features and semantics.
We described the design space in which these abstractions reside, their implicit or explicit use of stream and other non-MPI abstractions, their relationship to partitioned and persistent operations, and discussed their potential for added performance, how usable these abstractions are, and where functional and/or semantic gaps exist. Designs breaking backward compatibility with MPI were also duly noted (and cautioned). We also presented code examples of how some of these operate on the ubiquitous ping-pong use case (some utilize CUDA and others ROCM APIs).
The next steps to advance performance-portable programming with accelerator-based MPI+X programming should start with robust, inclusive conversations and design bake-offs among the various proposers, implementers, and users to identify a coherent path forward within the MPI community, inclusive of application developers. One or more proposed standard APIs should be competed and down-selected for MPI-5 (or an earlier MPI-4.x release). Significantly, the existing proposals/prototypes cited in this paper may not include all hardware vendors or systems currently available and certainly there may be other solutions forthcoming or that are not yet public (e.g., for FPGA-based accelerators or other GPU accelerator products). Yet, fitting additional and/or forthcoming proposals, prototypes, and designs into our taxonomy will provide a coherent basis to compare and contrast these with the eight presented here.
{credits}
6.0.1 Acknowledgements
This work was performed with partial support from the National Science Foundation under Grants Nos.  OAC-2103510, CCF-2405142 and CCF-2412182, the U.S. Department of Energy’s National Nuclear Security Administration (NNSA) under the Predictive Science Academic Alliance Program (PSAAP-III), Award DE-NA0003966, and Tennessee Technological University.
Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation, or the U.S. Department of Energy’s National Nuclear Security Administration.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.